byeeee:

What is the exact value of the expression the square root 72. − the square root of 8. + the square root of 128.? Simplify if possible. 8the square root of 2. 12the square root of 2. 8the square root of 3. 12the square root of 3.

Shadow:

\[\sqrt 72 - \sqrt 8 + \sqrt 128\] Correct?

byeeee:

yup

Shadow:

\[\sqrt 72 = \sqrt 8 \times 9 = \sqrt 2 \times 2 \times 2 \times 3 \times 3 = 2 \times 3 \sqrt 2 = 6 \sqrt 2\] \[\sqrt 8 = \sqrt 4 \times 2 = 2 \sqrt 2\] \[\sqrt 128 = \sqrt 4 \times 32 = \sqrt 4 \times 4 \times 4 \times 2 = 8 \sqrt 2\]

Shadow:

Does that make sense to you?

byeeee:

yes

Shadow:

\[6 \sqrt 2 - 2 \sqrt 2 + 8 \sqrt 2 = 12 \sqrt 2\] You basically only care about the number on the front. Since they are all multiplying by sqrt 2, you can treat them as like terms.

byeeee:

ohhhh!

byeeee:

thank you so much!